A Bibliography of Publications in Notices of the American Mathematical Society (1950–1959)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

15 August 2024 Version 1.00

Title word cross-reference

L [Tau59].

binary [SU59].

certain [SU59].

matrices [Tau59].

pairs [Tau59]. property [Tau59].

reaction [SU59].

systems [SU59].

weak [Tau59].

References

Stein:1959:CBR

[SU59] P. R. Stein and Stanisław M. Ulam. On certain binary reaction systems. Notices of the American Mathematical Society, 6(1):68–69, February 1959. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Taussky:1959:WPP

[Tau59] Olga Taussky. A weak property L for pairs of matrices. Notices of the American Mathematical Society, 6(1):69–70, February 1959. CODEN REFERENCES 2

AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).